Firms in a competitive market can sell as much as they like at a market price of $16. The cost function for each firm is TC = 50 + 4Q + 2Q^2 The associated marginal cost function is MC = 4 + 4Q and the point of minimum average cost is Q = 5. Using the profit maximizing rule, find the short-run profit at the profit maximizing quantity.
Answer #1

Answer

the profit is maximum at MC=P

4+4Q=16

4Q=12

Q=3

P=16

total revenue =P*Q=3*16=$48

total cost =50+4*3+2*3^2=80

Profit =TR-TC=48-80

=-$32

the short-run profit at the profit-maximizing quantity is -$32.